# Unlocking the Potential of Model Calibration in Federated Learning

### v2

@@ B.2 Additional Results of Calibration on FL
Calibration Method FedAvg McMahan et al. [2017] FedProx Sahu et al. [2018] Scaffold Karimireddy et al. [2019] FedDyn Acar et al. [2021] FedNova Wang et al. [2020b]
Uncal. 54.46 8.62 4.42 48.77 10.99 4.34 51.34 7.95 6.04 54.00 8.35 5.88 52.57 13.46 6.22
Focal Lin et al. [2017] 50.18 3.21 4.91 48.25 11.18 4.28 49.48 6.42 5.90 51.26 5.03 5.27 52.71 16.29 11.48
LS Müller et al. [2019] 52.13 2.92 4.75 47.23 6.99 5.28 51.94 7.27 6.79 54.12 4.09 3.68 52.22 12.66 6.02
BS Brier [1950] 47.18 5.21 5.02 42.16 5.21 3.36 48.09 7.42 5.69 53.62 9.23 6.33 50.25 16.54 6.40
MMCE Kumar et al. [2018] 52.62 8.25 3.95 48.76 6.91 4.93 51.90 8.02 6.02 53.99 8.53 6.17 51.50 14.95 9.13
FLSD Mukhoti et al. [2020] 48.31 5.21 5.65 48.21 9.25 4.45 49.03 8.53 7.43 54.69 10.53 6.70 50.58 14.67 9.42
MbLS Liu et al. [2021] 53.06 12.77 6.98 46.92 5.03 4.30 49.95 8.91 7.41 54.16 7.37 5.90 52.64 11.69 5.88
DCA Liang et al. [2020] 54.10 4.32 3.01 48.23 5.47 4.43 51.99 6.90 5.76 53.88 7.00 4.14 52.09 11.11 5.54
MDCA Hebbalaguppe et al. [2022b] 55.02 3.93 3.19 47.48 6.20 4.29 49.55 8.77 7.12 55.70 7.56 5.37 52.88 11.51 5.83
NUCFL (DCA+COS) 54.60 3.94 3.08 48.24 4.74 3.15 51.47 5.05 5.05 56.98 5.17 3.74 52.54 8.13 5.50
NUCFL (MDCA+COS) 55.25 3.74 2.89 48.54 4.37 3.13 51.62 5.38 5.16 56.03 5.03 4.06 52.68 9.54 5.11
NUCFL (DCA+L-CKA) 55.77 3.43 2.42 48.03 4.42 3.35 51.99 5.27 4.93 56.88 5.14 4.04 52.41 8.15 4.97
NUCFL (MDCA+L-CKA) 54.97 3.27 2.43 48.31 5.01 3.71 51.94 4.63 4.20 56.75 5.38 3.81 52.50 7.12 4.84
NUCFL (DCA+RBF-CKA) 55.60 3.39 2.25 48.88 4.70 3.71 52.17 5.14 5.09 56.77 5.27 3.90 52.20 7.19 4.18
NUCFL (MDCA+RBF-CKA) 55.97 3.46 2.21 48.32 4.84 3.74 52.04 5.06 4.66 56.88 5.44 3.99 52.45 7.36 5.02
@@
Table 13: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under non-IID ($\alpha=0.1$) scenario on the CIFAR-10 dataset. Values in boldface indicate the best calibration across all methods, achieved without compromising accuracy compared to the uncalibrated model.
@@
Calibration Method FedAvg McMahan et al. [2017] FedProx Sahu et al. [2018] Scaffold Karimireddy et al. [2019] FedDyn Acar et al. [2021] FedNova Wang et al. [2020b]
Uncal. 48.84 11.30 4.86 42.50 13.03 4.91 44.99 9.61 11.80 45.00 8.35 5.88 52.57 13.45 6.22
Focal Lin et al. [2017] 39.86 5.10 4.63 41.91 9.68 4.84 42.23 7.79 9.34 41.26 16.35 13.66 52.71 16.29 11.48
LS Müller et al. [2019] 43.13 7.08 5.82 41.91 6.72 5.38 43.30 9.03 10.88 45.12 7.35 4.37 52.22 12.68 6.02
BS Brier [1950] 31.80 14.48 6.67 41.75 6.98 6.96 41.20 7.42 9.48 44.62 9.23 6.33 50.25 16.54 6.40
MMCE Kumar et al. [2018] 43.72 7.59 3.91 42.96 8.52 3.93 42.50 7.69 9.72 43.39 9.78 8.69 51.50 14.95 9.13
FLSD Mukhoti et al. [2020] 34.83 5.22 6.42 42.22 6.01 5.39 42.65 9.15 10.69 44.63 9.60 6.80 50.58 14.67 9.42
MbLS Liu et al. [2021] 44.73 15.81 6.76 42.09 8.28 4.63 44.27 9.51 11.48 46.13 7.37 5.90 52.64 11.69 5.88
DCA Liang et al. [2020] 48.94 10.07 3.91 42.28 6.57 3.40 44.78 7.07 7.82 46.95 7.60 5.56 52.07 11.11 5.54
MDCA Hebbalaguppe et al. [2022b] 48.95 9.30 4.80 42.33 8.18 4.25 45.20 15.26 14.90 46.99 7.92 5.39 52.80 11.51 5.83
NUCFL (DCA+COS) 48.46 8.07 3.87 42.32 5.84 3.14 44.76 6.53 7.03 49.86 6.17 4.01 52.54 8.13 5.50
NUCFL (MDCA+COS) 48.67 8.53 4.17 43.20 5.01 2.75 44.37 6.52 7.14 50.31 6.03 4.06 52.68 9.54 5.11
NUCFL (DCA+L-CKA) 49.00 7.62 3.16 42.87 5.03 2.77 45.29 6.06 6.94 50.68 6.15 4.18 52.41 8.15 4.97
NUCFL (MDCA+L-CKA) 50.49 7.50 2.95 43.01 3.24 2.14 45.38 5.15 5.17 50.75 5.81 3.98 52.50 7.12 4.84
NUCFL (DCA+RBF-CKA) 49.50 7.76 3.09 42.90 4.90 2.61 45.99 6.14 6.48 49.67 6.27 4.31 52.19 7.19 4.18
NUCFL (MDCA+RBF-CKA) 49.32 7.72 3.07 43.08 3.44 2.13 45.69 5.39 5.59 49.89 6.18 4.28 52.45 7.36 5.02

### v3

@@ B.2 Additional Results of Calibration on FL
Calibration Method FedAvg (McMahan et al., 2017) FedProx (Sahu et al., 2018) Scaffold (Karimireddy et al., 2019)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 88.91 $\pm$ 0.53 5.58 $\pm$ 0.22 2.83 $\pm$ 0.35 88.67 $\pm$ 0.47 5.73 $\pm$ 0.16 2.79 $\pm$ 0.10 90.07 $\pm$ 0.62 6.67 $\pm$ 0.24 3.40 $\pm$ 0.17
Focal (Lin et al., 2017) 89.71 $\pm$ 0.48 6.89 $\pm$ 0.33 3.51 $\pm$ 0.22 89.06 $\pm$ 0.52 6.13 $\pm$ 0.17 3.06 $\pm$ 0.09 89.25 $\pm$ 0.57 7.71 $\pm$ 0.30 3.84 $\pm$ 0.17
LS (Müller et al., 2019) 88.89 $\pm$ 0.45 6.14 $\pm$ 0.17 3.06 $\pm$ 0.15 87.51 $\pm$ 0.49 6.49 $\pm$ 0.19 3.33 $\pm$ 0.26 90.11 $\pm$ 0.55 8.37 $\pm$ 0.66 3.41 $\pm$ 0.30
BS (Brier, 1950) 89.99 $\pm$ 0.52 8.26 $\pm$ 0.23 3.39 $\pm$ 0.17 90.51 $\pm$ 0.56 5.73 $\pm$ 0.58 2.77 $\pm$ 0.24 87.53 $\pm$ 0.47 6.23 $\pm$ 0.34 3.10 $\pm$ 0.22
MMCE (Kumar et al., 2018) 89.65 $\pm$ 0.92 8.13 $\pm$ 0.78 3.35 $\pm$ 0.44 88.66 $\pm$ 0.92 8.06 $\pm$ 0.40 3.30 $\pm$ 0.21 90.25 $\pm$ 1.08 7.15 $\pm$ 0.82 3.05 $\pm$ 0.39
FLSD (Mukhoti et al., 2020) 86.97 $\pm$ 0.67 5.41 $\pm$ 0.33 2.61 $\pm$ 0.23 89.73 $\pm$ 0.71 9.09 $\pm$ 0.32 3.49 $\pm$ 0.19 90.91 $\pm$ 0.71 8.46 $\pm$ 0.30 3.36 $\pm$ 0.22
MbLS (Liu et al., 2021) 86.93 $\pm$ 1.09 7.39 $\pm$ 0.77 3.59 $\pm$ 0.39 89.43 $\pm$ 0.98 7.62 $\pm$ 0.41 3.82 $\pm$ 0.32 89.05 $\pm$ 0.77 8.24 $\pm$ 0.21 3.39 $\pm$ 0.20
DCA (Liang et al., 2020) 89.04 $\pm$ 0.57 4.79 $\pm$ 0.22 1.98 $\pm$ 0.14 89.90 $\pm$ 0.61 4.91 $\pm$ 0.23 2.51 $\pm$ 0.15 90.16 $\pm$ 0.64 6.08 $\pm$ 0.24 3.00 $\pm$ 0.17
NUCFL (DCA+COS) 90.01 $\pm$ 0.72 4.50 $\pm$ 0.34 1.77 $\pm$ 0.19 88.86 $\pm$ 0.88 4.62 $\pm$ 0.39 1.72 $\pm$ 0.25 90.80 $\pm$ 0.71 5.86 $\pm$ 0.41 2.69 $\pm$ 0.29
NUCFL (DCA+L-CKA) 89.89 $\pm$ 0.53 4.49 $\pm$ 0.19 1.77 $\pm$ 0.12 88.99 $\pm$ 0.57 4.56 $\pm$ 0.20 1.69 $\pm$ 0.11 90.23 $\pm$ 0.60 5.61 $\pm$ 0.39 2.66 $\pm$ 0.20
NUCFL (DCA+RBF-CKA) 89.07 $\pm$ 0.61 4.38 $\pm$ 0.25 1.71 $\pm$ 0.14 89.61 $\pm$ 0.58 4.77 $\pm$ 0.22 1.88 $\pm$ 0.12 90.31 $\pm$ 0.61 5.70 $\pm$ 0.23 2.68 $\pm$ 0.14
MDCA (Hebbalaguppe et al., 2022b) 89.18 $\pm$ 0.37 4.88 $\pm$ 0.22 2.34 $\pm$ 0.13 $8$$\pm$ 0.58 4.87 $\pm$ 0.24 2.50 $\pm$ 0.15 91.03 $\pm$ 0.60 5.95 $\pm$ 0.26 2.98 $\pm$ 0.16
NUCFL (MDCA+COS) 89.03 $\pm$ 1.00 4.44 $\pm$ 0.39 1.78 $\pm$ 0.28 88.74 $\pm$ 0.91 4.77 $\pm$ 0.67 1.90 $\pm$ 0.32 90.52 $\pm$ 0.55 5.79 $\pm$ 0.33 2.67 $\pm$ 0.18
NUCFL (MDCA+L-CKA) 89.11 $\pm$ 0.78 4.39 $\pm$ 0.61 1.73 $\pm$ 0.30 88.96 $\pm$ 0.51 4.61 $\pm$ 0.22 1.72 $\pm$ 0.23 90.47 $\pm$ 0.54 5.67 $\pm$ 0.23 2.66 $\pm$ 0.14
NUCFL (MDCA+RBF-CKA) 88.99 $\pm$ 0.48 4.87 $\pm$ 0.21 2.35 $\pm$ 0.15 89.91 $\pm$ 0.52 4.58 $\pm$ 0.22 1.69 $\pm$ 0.13 90.84 $\pm$ 0.53 5.67 $\pm$ 0.24 2.65 $\pm$ 0.14
Calibration Method FedDyn (Acar et al., 2021) FedNova (Wang et al., 2020b)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 89.03 $\pm$ 0.58 6.02 $\pm$ 0.23 3.00 $\pm$ 0.17 89.56 $\pm$ 0.51 5.88 $\pm$ 0.22 2.76 $\pm$ 0.19
Focal (Lin et al., 2017) 89.13 $\pm$ 0.30 8.48 $\pm$ 0.22 3.45 $\pm$ 0.20 89.42 $\pm$ 0.49 6.12 $\pm$ 0.30 3.07 $\pm$ 0.21
LS (Müller et al., 2019) 88.34 $\pm$ 0.53 9.03 $\pm$ 0.44 3.50 $\pm$ 0.23 88.63 $\pm$ 0.50 7.57 $\pm$ 0.25 3.20 $\pm$ 0.10
BS (Brier, 1950) 88.90 $\pm$ 0.71 7.38 $\pm$ 0.33 3.59 $\pm$ 0.14 88.94 $\pm$ 0.48 7.43 $\pm$ 0.25 3.61 $\pm$ 0.14
MMCE (Kumar et al., 2018) 88.80 $\pm$ 1.57 7.90 $\pm$ 1.01 3.33 $\pm$ 0.49 88.94 $\pm$ 1.05 7.11 $\pm$ 0.77 3.04 $\pm$ 0.39
FLSD (Mukhoti et al., 2020) 88.90 $\pm$ 0.57 7.18 $\pm$ 0.24 3.05 $\pm$ 0.13 87.45 $\pm$ 1.04 6.09 $\pm$ 0.88 2.99 $\pm$ 0.39
MbLS (Liu et al., 2021) 88.10 $\pm$ 0.80 7.33 $\pm$ 0.37 3.59 $\pm$ 0.11 86.74 $\pm$ 0.91 5.14 $\pm$ 0.41 2.60 $\pm$ 0.29
DCA (Liang et al., 2020) 90.20 $\pm$ 0.66 5.72 $\pm$ 0.25 2.77 $\pm$ 0.16 90.11 $\pm$ 0.59 5.00 $\pm$ 0.22 2.45 $\pm$ 0.14
NUCFL (DCA+COS) 90.10 $\pm$ 0.56 5.33 $\pm$ 0.22 2.58 $\pm$ 0.28 90.96 $\pm$ 0.54 4.79 $\pm$ 0.40 2.00 $\pm$ 0.39
NUCFL (DCA+L-CKA) 89.37 $\pm$ 0.77 5.17 $\pm$ 0.23 2.55 $\pm$ 0.14 90.57 $\pm$ 1.39 4.62 $\pm$ 0.89 1.97 $\pm$ 0.30
NUCFL (DCA+RBF-CKA) 89.59 $\pm$ 0.63 5.40 $\pm$ 0.39 2.61 $\pm$ 0.28 90.19 $\pm$ 1.01 4.80 $\pm$ 0.72 1.99 $\pm$ 0.31
MDCA (Hebbalaguppe et al., 2022b) 90.10 $\pm$ 0.62 5.81 $\pm$ 0.25 2.79 $\pm$ 0.14 90.03 $\pm$ 0.57 5.27 $\pm$ 0.23 2.50 $\pm$ 0.12
NUCFL (MDCA+COS) 90.03 $\pm$ 0.58 5.45 $\pm$ 0.49 2.63 $\pm$ 0.20 90.36 $\pm$ 0.71 5.15 $\pm$ 0.69 2.50 $\pm$ 0.30
NUCFL (MDCA+L-CKA) 90.57 $\pm$ 1.56 5.66 $\pm$ 1.24 2.69 $\pm$ 0.70 89.62 $\pm$ 0.33 4.93 $\pm$ 0.22 2.05 $\pm$ 0.19
NUCFL (MDCA+RBF-CKA) 89.30 $\pm$ 0.71 5.73 $\pm$ 0.38 2.70 $\pm$ 0.29 89.66 $\pm$ 0.53 4.69 $\pm$ 0.31 1.99 $\pm$ 0.09
@@
Table 13: Accuracy (%), calibration measures ECE (%), and SCE (%) of various FL algorithms with different calibration methods under non-IID ($\alpha=0.1$) scenario on the FEMNIST dataset. Underlined values indicate the best calibration across all methods.
@@
Calibration Method FedAvg (McMahan et al., 2017) FedProx (Sahu et al., 2018) Scaffold (Karimireddy et al., 2019)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 86.98 $\pm$ 0.92 8.78 $\pm$ 0.28 3.49 $\pm$ 0.18 87.57 $\pm$ 1.01 7.03 $\pm$ 0.32 3.68 $\pm$ 0.17 87.17 $\pm$ 0.95 9.73 $\pm$ 0.35 3.70 $\pm$ 0.19
Focal (Lin et al., 2017) 84.71 $\pm$ 1.10 10.09 $\pm$ 0.77 3.55 $\pm$ 0.28 86.06 $\pm$ 1.26 9.33 $\pm$ 0.77 3.42 $\pm$ 0.69 87.25 $\pm$ 1.00 11.99 $\pm$ 0.36 4.35 $\pm$ 0.25
LS (Müller et al., 2019) 84.89 $\pm$ 0.91 8.34 $\pm$ 0.39 3.35 $\pm$ 0.22 87.01 $\pm$ 0.97 8.69 $\pm$ 0.33 3.49 $\pm$ 0.29 86.91 $\pm$ 0.92 9.67 $\pm$ 0.33 3.71 $\pm$ 0.11
BS (Brier, 1950) 85.99 $\pm$ 0.95 9.46 $\pm$ 0.34 3.44 $\pm$ 0.20 86.51 $\pm$ 1.03 8.93 $\pm$ 0.88 3.44 $\pm$ 0.37 86.53 $\pm$ 0.89 9.43 $\pm$ 0.27 3.69 $\pm$ 0.16
MMCE (Kumar et al., 2018) 86.65 $\pm$ 1.01 9.33 $\pm$ 0.47 3.43 $\pm$ 0.39 85.66 $\pm$ 0.88 6.26 $\pm$ 0.57 3.08 $\pm$ 0.19 87.25 $\pm$ 0.94 13.35 $\pm$ 0.32 4.52 $\pm$ 0.19
FLSD (Mukhoti et al., 2020) 85.97 $\pm$ 0.88 9.61 $\pm$ 0.35 3.47 $\pm$ 0.23 86.73 $\pm$ 0.91 9.29 $\pm$ 0.29 3.42 $\pm$ 0.17 87.91 $\pm$ 1.79 14.66 $\pm$ 1.46 5.07 $\pm$ 0.72
MbLS (Liu et al., 2021) 85.93 $\pm$ 0.77 9.59 $\pm$ 0.66 3.51 $\pm$ 0.28 86.43 $\pm$ 1.22 8.82 $\pm$ 0.61 3.47 $\pm$ 0.32 86.05 $\pm$ 1.33 10.44 $\pm$ 0.72 3.81 $\pm$ 0.38
DCA (Liang et al., 2020) 87.14 $\pm$ 1.04 8.39 $\pm$ 0.24 3.41 $\pm$ 0.11 87.40 $\pm$ 1.22 6.13 $\pm$ 0.49 3.06 $\pm$ 0.31 88.05 $\pm$ 1.12 9.08 $\pm$ 0.44 3.52 $\pm$ 0.31
NUCFL (DCA+COS) 87.07 $\pm$ 1.02 8.02 $\pm$ 0.26 3.30 $\pm$ 0.17 88.16 $\pm$ 1.25 5.96 $\pm$ 0.33 3.01 $\pm$ 0.21 87.80 $\pm$ 1.07 8.79 $\pm$ 0.61 3.49 $\pm$ 0.44
NUCFL (DCA+L-CKA) 86.89 $\pm$ 1.24 7.84 $\pm$ 0.71 3.27 $\pm$ 0.33 87.56 $\pm$ 1.04 5.06 $\pm$ 0.29 2.54 $\pm$ 0.17 87.23 $\pm$ 1.34 8.51 $\pm$ 0.72 3.45 $\pm$ 0.31
NUCFL (DCA+RBF-CKA) 86.93 $\pm$ 1.05 7.80 $\pm$ 0.49 3.25 $\pm$ 0.22 88.11 $\pm$ 1.21 5.09 $\pm$ 0.46 2.53 $\pm$ 0.25 88.31 $\pm$ 1.03 8.47 $\pm$ 0.66 3.46 $\pm$ 0.29
MDCA (Hebbalaguppe et al., 2022b) 87.07 $\pm$ 1.12 8.22 $\pm$ 0.58 3.34 $\pm$ 0.35 89.06 $\pm$ 1.15 6.28 $\pm$ 0.38 3.08 $\pm$ 0.11 88.23 $\pm$ 1.18 8.99 $\pm$ 0.43 3.51 $\pm$ 0.27
NUCFL (MDCA+COS) 87.03 $\pm$ 1.10 8.20 $\pm$ 0.41 3.36 $\pm$ 0.27 88.74 $\pm$ 1.14 5.71 $\pm$ 0.38 2.79 $\pm$ 0.26 87.52 $\pm$ 1.33 8.59 $\pm$ 0.87 3.45 $\pm$ 0.41
NUCFL (MDCA+L-CKA) 87.11 $\pm$ 1.33 7.99 $\pm$ 0.92 3.30 $\pm$ 0.61 87.96 $\pm$ 1.42 5.34 $\pm$ 0.91 2.70 $\pm$ 0.33 87.47 $\pm$ 1.00 8.61 $\pm$ 0.76 3.46 $\pm$ 0.41
NUCFL (MDCA+RBF-CKA) 87.39 $\pm$ 1.31 7.87 $\pm$ 0.61 3.25 $\pm$ 0.44 88.01 $\pm$ 1.14 5.29 $\pm$ 0.41 2.67 $\pm$ 0.29 87.84 $\pm$ 1.31 8.65 $\pm$ 0.41 3.47 $\pm$ 0.28
Calibration Method FedDyn (Acar et al., 2021) FedNova (Wang et al., 2020b)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 88.03 $\pm$ 1.08 10.51 $\pm$ 0.33 3.66 $\pm$ 0.21 87.56 $\pm$ 1.02 9.08 $\pm$ 0.30 3.50 $\pm$ 0.18
Focal (Lin et al., 2017) 87.13 $\pm$ 1.03 11.28 $\pm$ 0.77 3.69 $\pm$ 0.39 88.00 $\pm$ 0.94 12.92 $\pm$ 0.49 4.47 $\pm$ 0.35
LS (Müller et al., 2019) 88.34 $\pm$ 1.44 15.23 $\pm$ 1.31 5.16 $\pm$ 0.72 88.63 $\pm$ 0.98 10.77 $\pm$ 0.33 3.69 $\pm$ 0.25
BS (Brier, 1950) 87.90 $\pm$ 0.97 12.58 $\pm$ 0.54 4.45 $\pm$ 0.21 86.94 $\pm$ 0.94 10.13 $\pm$ 0.28 3.64 $\pm$ 0.15
MMCE (Kumar et al., 2018) 87.80 $\pm$ 1.19 15.10 $\pm$ 0.48 5.13 $\pm$ 0.22 87.11 $\pm$ 1.03 13.31 $\pm$ 0.73 4.52 $\pm$ 0.49
FLSD (Mukhoti et al., 2020) 86.90 $\pm$ 0.94 9.07 $\pm$ 0.53 3.52 $\pm$ 0.34 86.45 $\pm$ 1.12 10.29 $\pm$ 0.49 3.65 $\pm$ 0.30
MbLS (Liu et al., 2021) 87.16 $\pm$ 1.07 15.53 $\pm$ 0.61 5.17 $\pm$ 0.32 85.74 $\pm$ 0.82 14.84 $\pm$ 0.30 5.06 $\pm$ 0.29
DCA (Liang et al., 2020) 88.70 $\pm$ 1.11 8.92 $\pm$ 0.41 3.45 $\pm$ 0.19 89.06 $\pm$ 1.08 8.60 $\pm$ 0.42 3.47 $\pm$ 0.29
NUCFL (DCA+COS) 88.10 $\pm$ 1.09 8.53 $\pm$ 0.41 3.41 $\pm$ 0.30 87.96 $\pm$ 1.06 8.49 $\pm$ 0.30 3.41 $\pm$ 0.18
NUCFL (DCA+L-CKA) 88.37 $\pm$ 1.09 7.87 $\pm$ 0.24 3.26 $\pm$ 0.11 87.56 $\pm$ 1.03 8.38 $\pm$ 0.47 3.39 $\pm$ 0.22
NUCFL (DCA+RBF-CKA) 88.59 $\pm$ 1.10 8.60 $\pm$ 0.57 3.41 $\pm$ 0.44 87.34 $\pm$ 1.22 8.30 $\pm$ 0.17 3.37 $\pm$ 0.09
MDCA (Hebbalaguppe et al., 2022b) 88.10 $\pm$ 1.16 9.18 $\pm$ 0.44 3.54 $\pm$ 0.31 87.73 $\pm$ 1.33 8.67 $\pm$ 0.69 3.47 $\pm$ 0.42
NUCFL (MDCA+COS) 88.11 $\pm$ 1.00 8.75 $\pm$ 0.49 3.49 $\pm$ 0.32 88.00 $\pm$ 1.24 7.55 $\pm$ 0.61 3.27 $\pm$ 0.35
NUCFL (MDCA+L-CKA) 88.07 $\pm$ 1.13 8.66 $\pm$ 0.52 3.42 $\pm$ 0.29 87.59 $\pm$ 1.23 7.51 $\pm$ 0.61 3.28 $\pm$ 0.39
NUCFL (MDCA+RBF-CKA) 88.30 $\pm$ 1.29 8.83 $\pm$ 0.44 3.50 $\pm$ 0.31 87.96 $\pm$ 1.24 7.80 $\pm$ 0.39 3.33 $\pm$ 0.22